40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4816     Entry time: Tue Jun 14 12:23:44 2011
Author: Jamie, Joe 
Type: Update 
Category: CDS 
Subject: WE ARE ALL GREEN! LSC back up and running in new configuration. 

After moving the c1lsc computer to 1X4, then connecting c1lsc to it's IO chassis in 1Y3 by a fiber PCIe extension cable, everything is back up and running and the status screen is all green.  c1lsc is now directly connected to c1sus via a short copper Dolphin cable.

After lunch we will do some more extensive testing of the system to make sure everything is working as expected.

ELOG V3.1.3-